If Ali travels to his school at the speed of 5 mph, he is late by 20 min. If he travels at the speed of 6 mph, he is late by 4 min. What is the distance between Ali's home and his school ( in miles) ?

Thanks.

Answer by richwmiller(9135)

r*t=d
5*(t+20/60)=6*(t+4/60)
5*(t+1/3)=6*(t+1/15)
5t+5/3=6t+6/15
5t+25/15=6t+6/15
t=19/15
5*(19/15+20/60)=8 miles
and
6*(19/15+4/60)=8 miles